# A Survey of Reinforcement Learning For Economics

### 09 Mar 2026 v1

@@ 4.5.1 Actor-Critic Architecture and Two-Timescale Convergence
Convergence requires the critic to learn faster than the actor:
@@
Convergence requires a structural condition on the critic. The critic’s feature vectors must span the actor’s score functions $\nabla_{\omega}\log\pi_{\omega}(a|s)$, so that the critic’s approximation error lies orthogonal to the policy gradient direction. Under this compatibility condition, the critic’s projection error does not bias the actor’s gradient estimates. [^68] The compatible function approximation theorem first appears in Sutton et al. (2000) and is the key structural requirement in Konda and Tsitsiklis (2000). It constrains the critic architecture to be “compatible” with the actor parameterization — the same condition that makes the natural policy gradient equal to the critic’s weight vector in Kakade (2002).
@@ 4.5.2 Entropy Regularization and Soft Actor-Critic
where $\mathcal{H}(\pi)=-\sum_{a}\pi(a)\log\pi(a)$ is the entropy and $\tau>0$ is the temperature parameter. Geist et al. (2019) later provided the unifying theoretical framework, showing that entropy regularization converts the Bellman optimality operator’s non-smooth hard max into a smooth log-sum-exp, and that the resulting soft Bellman operator remains a $\gamma$-contraction, preserving the convergence guarantees of standard dynamic programming. [^70] The optimal policy under entropy regularization is $\pi^{*}(a|s)\propto\exp(Q^{*}(s,a)/\tau)$, which is precisely the McFadden (1974b) logit choice probability with systematic utility $Q^{*}(s,a)$ and scale parameter $\tau$. The entropy-regularized value function is the log-sum-exp of Q-values, corresponding to the inclusive value (log-sum) operator in nested logit models. This equivalence between the soft-control framework and dynamic discrete choice models with EV1 taste shocks is developed formally in Rust and Rawat (2026), Appendix A.
@@ 4.6 Fundamental Tradeoffs
The choice between methods involves distinct tradeoffs rooted in DP structure. Value-based methods target $Q^{*}$ via the Bellman contraction (Szepesvári, 2010). In the tabular case convergence is guaranteed, but function approximation introduces the deadly triad. Policy-based methods optimize $\pi_{\theta}$ directly. Modern theory (Agarwal et al., 2021) establishes global convergence for softmax policies, with high variance as the practical weakness rather than local traps. Actor-critic methods combine both (Konda and Tsitsiklis, 2000), using the critic for low-variance value estimates while the actor inherits policy gradient’s global convergence. Each family traces to DP foundations.
Four additional trade-offs pervade reinforcement learning. First, exploration versus exploitation: should the agent act on its current best estimate or gather information to improve future decisions? Lai and Robbins (1985) establish the fundamental lower bound: any consistent policy must incur regret at least logarithmic in the number of periods. Naive exploration ($\varepsilon$-greedy) requires samples exponential in the horizon; strategic exploration (UCB, optimism in the face of uncertainty) reduces this to polynomial (Auer et al., 2002), formalizing the value of targeted experimentation. [^75] The exploration-exploitation tradeoff is the subject of the bandits chapter, where the multi-armed bandit framework provides the sharpest analysis. In full MDPs, exploration is harder because the agent must learn not just reward distributions but also transition dynamics, compounding the information requirement.
@@ 4.7 Conclusion
The central insight is that RL algorithms are not mysterious. They are asymptotic approximations to classical dynamic programming operators, justified by the mathematics of contractions, stochastic approximation, and gradient domination. Value iteration becomes Q-learning (Watkins and Dayan, 1992; Tsitsiklis, 1994) when expectations are replaced by single samples. Policy iteration becomes the natural policy gradient (Kakade, 2001; Agarwal et al., 2021) when the greedy improvement step is approximated by gradient ascent, and NPG recovers PI exactly in the tabular case. The stochastic approximation framework, from the foundational work of Robbins (1952) through the ODE method of Borkar and Meyn (2000), guarantees that under appropriate step-size conditions, noisy iterates converge to the same fixed points as their deterministic counterparts. Reinforcement learning is not a departure from dynamic programming but an extension of it. Tabular RL and RL with linear function approximation is well understood. However, deep RL is not as well understood. It seems to work for some cases but fails in others, and we do not have a firm handle on why. But in some sense, we do know that when it works it works due to the underlying Bellman principle. For these reasons convergence in RL remains and open issue and successes in RL remain brittle and case-specific.

### 15 Mar 2026 v2

@@ 5.5.1 Actor-Critic Architecture and Two-Timescale Convergence
Convergence requires the critic to learn faster than the actor.
@@
Convergence requires a structural condition on the critic. The critic’s feature vectors must span the actor’s score functions $\nabla_{\omega}\log\pi_{\omega}(a|s)$, so that the critic’s approximation error lies orthogonal to the policy gradient direction. Under this compatibility condition, the critic’s projection error does not bias the actor’s gradient estimates. [^84] The compatible function approximation theorem first appears in Sutton et al. (2000) and is the key structural requirement in Konda and Tsitsiklis (2000). It constrains the critic architecture to be “compatible” with the actor parameterization, the same condition that makes the natural policy gradient equal to the critic’s weight vector in Kakade (2002).
@@ 5.5.2 Entropy Regularization and Soft Actor-Critic
where $\mathcal{H}(\pi)=-\sum_{a}\pi(a)\log\pi(a)$ is the entropy and $\tau>0$ is the temperature parameter. Geist et al. (2019) later provided the unifying theoretical framework, showing that entropy regularization converts the Bellman optimality operator’s non-smooth hard max into a smooth log-sum-exp, and that the resulting soft Bellman operator remains a $\gamma$-contraction, preserving the convergence guarantees of standard dynamic programming. [^86] The optimal policy under entropy regularization is $\pi^{*}(a|s)\propto\exp(Q^{*}(s,a)/\tau)$, which is precisely the McFadden (1974) logit choice probability with systematic utility $Q^{*}(s,a)$ and scale parameter $\tau$. The entropy-regularized value function is the log-sum-exp of Q-values, corresponding to the inclusive value (log-sum) operator in nested logit models. This equivalence between the soft-control framework and dynamic discrete choice models with EV1 taste shocks is developed formally in Rust and Rawat (2026), Appendix A.
@@ 5.6 Fundamental Tradeoffs
The choice between methods involves distinct tradeoffs rooted in DP structure. Value-based methods target $Q^{*}$ via the Bellman contraction (Szepesvári, 2010). In the tabular case convergence is guaranteed, but function approximation introduces the deadly triad. Policy-based methods optimize $\pi_{\theta}$ directly. Modern theory (Agarwal et al., 2021a) establishes global convergence for softmax policies, with high variance as the practical weakness rather than local traps. Actor-critic methods combine both (Konda and Tsitsiklis, 2000), using the critic for low-variance value estimates while the actor inherits policy gradient’s global convergence. Each family traces to DP foundations.
Four additional trade-offs pervade reinforcement learning. First, exploration versus exploitation: should the agent act on its current best estimate or gather information to improve future decisions? Lai and Robbins (1985) establish the fundamental lower bound: any consistent policy must incur regret at least logarithmic in the number of periods. Naive exploration ($\varepsilon$-greedy) requires samples exponential in the horizon; strategic exploration (UCB, optimism in the face of uncertainty) reduces this to polynomial (Auer et al., 2002a), formalizing the value of targeted experimentation. [^91] The exploration-exploitation tradeoff is the subject of the bandits chapter, where the multi-armed bandit framework provides the sharpest analysis. In full MDPs, exploration is harder because the agent must learn not just reward distributions but also transition dynamics, compounding the information requirement.
@@ 5.7 Conclusion
The central insight is that RL algorithms are not mysterious. They are asymptotic approximations to classical dynamic programming operators, justified by the mathematics of contractions, stochastic approximation, and gradient domination. Value iteration becomes Q-learning (Watkins and Dayan, 1992; Tsitsiklis, 1994) when expectations are replaced by single samples. Policy iteration becomes the natural policy gradient (Kakade, 2001; Agarwal et al., 2021a) when the greedy improvement step is approximated by gradient ascent, and NPG recovers PI exactly in the tabular case. The stochastic approximation framework, from the foundational work of Robbins (1952) through the ODE method of Borkar and Meyn (2000), guarantees that under appropriate step-size conditions, noisy iterates converge to the same fixed points as their deterministic counterparts. Reinforcement learning is not a departure from dynamic programming but an extension of it. Tabular RL and RL with linear function approximation rest on solid theoretical foundations. Deep RL lacks comparable guarantees: convergence remains an open problem, and empirical successes remain case-specific.
6 The Empirics of Deep Reinforcement Learning
I review the empirical pathologies of deep reinforcement learning, their causes, and the diagnostic tools that expose them.
6.1 The Moving Target Problem
In supervised learning, the loss function is a fixed function of the training data and model parameters. Deep reinforcement learning does not enjoy this property. Each gradient step moves both the current value estimates and the targets simultaneously, creating a “nonstationary” optimization landscape. The target network heuristic introduced by Mnih et al. (2015) slows target drift by periodically freezing a copy of the network, but does not eliminate it. Therefore Bellman residual is a poor proxy for the accuracy of the value function.
Fujimoto et al. (2022) formalize this observation. Let $Q^{\pi}$ denote the true value function for policy $\pi$, let $\Delta(s,a)=Q(s,a)-Q^{\pi}(s,a)$ denote the value error, and let $\varepsilon(s,a)=Q(s,a)-(r+\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[Q(s^{\prime},a^{\prime})])$ denote the Bellman error. Substituting the definition of $Q^{\pi}$ yields the key identity: $\varepsilon(s,a)=\Delta(s,a)-\gamma\mathbb{E}_{s^{\prime},a^{\prime}}[\Delta(s^{\prime},a^{\prime})]$. The Bellman error is a difference of value errors at consecutive states, not the value error itself. If the errors $\Delta(s,a)$ and $\Delta(s^{\prime},a^{\prime})$ are correlated across time—the network is wrong in the same direction at successive state-action pairs—they cancel in the difference and the Bellman error is small regardless of how large the individual errors are. [^93] In the extreme case, shifting all Q-values by the constant $c/(1-\gamma)$ leaves the Bellman error unchanged at zero while increasing value error by $c/(1-\gamma)$.
The second failure mode is specific to finite datasets: Fujimoto et al. (2022, Corollary 1) show that over an incomplete dataset, zero Bellman error is consistent with arbitrarily large value error, because the network can fit unobserved successor pairs to whatever values make the observed residuals vanish. [^94] The Bellman equation uniquely identifies $Q^{\pi}$ when enforced over the entire MDP, but over an incomplete dataset it admits infinitely many solutions. Whenever a transition $(s^{\prime},a^{\prime})$ that is reachable from the dataset is not itself in the dataset, the network is free to set $Q(s^{\prime},a^{\prime})$ at unobserved pairs to whatever value makes the residual vanish on the observed ones, unconstrained by any loss. A network can thus reach near-zero training loss while the value function remains arbitrarily inaccurate over the full state space.
The dual failure mode appears in policy gradient methods: Ilyas et al. (2020) find that even when PPO’s surrogate objective improves monotonically, episode return can plateau or decline, because the surrogate gradient is poorly aligned with the gradient of the true return. [^95] Ilyas et al. (2020) examine the surrogate objective used in Proximal Policy Optimization (Schulman et al., 2017). The PPO clipping mechanism is designed to keep policy updates within a trust region by bounding the probability ratio $\pi_{\theta}(a|s)/\pi_{\theta_{\text{old}}}(a|s)$. The gradient of the surrogate is poorly aligned with the gradient of the true return, particularly in later training phases where the policy has diverged from the behavior policy used to collect the replay data. The loss metric that practitioners monitor throughout training is measuring something other than what they care about.
6.2 The Reproducibility Crisis and Sensitivity to Random Seeds
Henderson et al. (2018) trained five leading policy gradient algorithms (PPO, TRPO, DDPG, TD3, SAC) on six MuJoCo benchmark environments, holding all hyperparameters fixed and varying only the random seed. The resulting learning curves from different seeds were non-overlapping: a seed that performed well under one algorithm performed comparably to a different algorithm’s best seeds, making cross-algorithm comparison unreliable. [^96] Differences as large as 2,000 points in final episode return arose from seed variation alone. Agarwal et al. (2021b) quantify the damage: comparing point estimates from 5 runs per task on Atari 100k yields Type I error exceeding 50%, meaning a random noise injection appears beneficial in half of all comparisons.
Agarwal et al. (2021b) propose the interquartile mean (IQM) as a replacement for mean and median when comparing algorithms. The IQM discards the top and bottom 25% of runs before averaging, reducing sensitivity to outlier seeds. Using these tools and stratified bootstrap confidence intervals, Agarwal et al. (2021b) find that several widely-cited algorithmic improvements on Atari 100k vanish or reverse when statistical uncertainty is accounted for. [^97] They also introduce performance profiles, which plot the fraction of tasks and seeds where an algorithm achieves performance above a threshold $\tau$, as $\tau$ varies from 0 to the maximum. Performance profiles reveal the full shape of the score distribution rather than collapsing it to a single statistic. [^98] The fragility extends to hyperparameters. Eimer et al. (2023) conduct a systematic study of hyperparameter sensitivity across 6 algorithms and 17 environments, finding that default hyperparameters from published papers perform competitively in the specific environments used in those papers but generalize poorly across environments. Patterson et al. (2024) synthesize these findings into an empirical design handbook, recommending at minimum 10 seeds per configuration, IQM-based comparisons, and preregistration of hyperparameter search protocols.
6.3 Value Overestimation and Spikes
Q-learning uses the Bellman optimality update
$Q(s,a)\leftarrow r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime}),$ (58)
where the maximum is taken over the estimated Q-values of all actions at the successor state. Thrun and Schwartz (1993) identify a positive bias intrinsic to this update: if the Q-value estimates contain noise with mean zero, the maximum over noisy estimates is biased upward by Jensen’s inequality. An agent that uses a single network for both action selection ($\arg\max$) and value estimation ($\max Q$) systematically overestimates the values of every state it visits, biasing the Bellman target upward at every update step. The bias compounds through bootstrapping: overestimated targets produce overestimated updates, which produce further overestimated targets.
van Hasselt (2010) propose double Q-learning as a remedy: maintain two independent Q-networks $Q_{A}$ and $Q_{B}$. Use $Q_{A}$ to select the greedy action at $s^{\prime}$, but use $Q_{B}$ to evaluate that action. Because the two networks are trained on different data, their errors are approximately independent, and the positive bias largely cancels. van Hasselt et al. (2016b) implement this as Double DQN, using the online network for action selection and the periodically-frozen target network for evaluation. On 49 Atari games, Double DQN reduces overestimation by a factor of 3–5 and improves median performance by 20% relative to DQN.
Fujimoto et al. (2018) observe that Double DQN’s correction is incomplete in continuous-action settings, where the target network and online network remain correlated through shared updates. They propose Clipped Double Q-learning: compute two Q-value estimates $Q_{1},Q_{2}$ with separate networks trained on the same data, and use $y=r+\gamma\min(Q_{1}(s^{\prime},a^{\prime}),Q_{2}(s^{\prime},a^{\prime}))$ as the Bellman target. The minimum operator introduces pessimistic underestimation, which is conservative but avoids the explosive positive bias. [^99] Ciosek et al. (2019) note that clipped double Q can cause excessive pessimism under high uncertainty, proposing optimistic actor-critic as a counterweight.
Figure 11: Overestimation bias from Jensen’s inequality with $n=2$ actions. Blue: density of individual $\hat{Q}(s,a_{i})\sim\mathcal{N}(\mu,\sigma^{2})$. Red: density of $\max_{i}\hat{Q}(s,a_{i})$, shifted right by $\sigma/\sqrt{\pi}\approx 0.56$. The shaded gap is the bias. With $n=100$ actions the bias exceeds $2.5\sigma$.
DQN prevents outright divergence, but van Hasselt et al. (2018) find that soft divergence, defined as a temporary spike in value estimates by more than 10% followed by recovery, occurs in the majority of DQN training runs across 57 Atari games.100100100Larger networks diverge more frequently than smaller ones, counter to the usual intuition that more expressive models should generalize better. The deadly triad does not announce itself as a training failure: the value estimates may spike and recover, leaving no visible trace in the loss curve while corrupting the policy.
Kumar et al. (2021) describe a continuous manifestation of the deadly triad: implicit under-parameterization.101101101Neural networks trained with bootstrapped TD objectives progressively lose their effective rank, with fewer and fewer neurons contributing distinct directions in the representation. This rank collapse is silent (training loss continues to decrease) but the network’s ability to represent new information degrades over time, a form of capacity loss distinct from but related to plasticity loss.
6.4 Plasticity Loss and Primacy Bias
A network that trains well at step $t=100$ may be incapable of learning at step t=100,000t=100{,}000, even if the data quality at the later step is higher. Lyle et al. (2022) call this capacity loss: the network’s ability to update its own weights degrades progressively during training, measured by the fraction of effective parameters and the network’s ability to fit new random labels. Lyle et al. (2023) extend this to plasticity loss, identifying dead ReLU neurons, weight norm growth, and feature rank collapse as three distinct mechanisms, not all of which co-occur.
Nikishin et al. (2022) identify primacy bias as a specific cause. Because the replay buffer is filled incrementally, early transitions are oversampled relative to later ones, and the network over-fits to early environment transitions, corrupting representations throughout the remainder of training.102102102Nikishin et al. (2022) show that 100 initial “priming” steps of excessive gradient updates degrade a SAC agent’s performance for hundreds of thousands of subsequent steps. Sokar et al. (2023) measure the fraction of dormant neurons (units with near-zero activation across the replay buffer) accumulating monotonically during training. Dohare et al. (2024) find that standard deep networks lose all plasticity within a few million gradient steps in continual learning tasks.
The proposed remedies fall into three categories: periodic resets, continual backpropagation, and architectural interventions.103103103Periodic resets reinitialize the last layers of the network while retaining the replay buffer, allowing the agent to forget overfit representations without discarding experience (Nikishin et al., 2022; D’Oro et al., 2023). Continual backpropagation replaces neurons with near-zero utility at each gradient step rather than waiting for a global reset (Dohare et al., 2024). Architectural interventions—layer normalization (Lyle et al., 2025), orthogonal initialization, spectral normalization—reduce the rate at which plasticity is lost by stabilizing gradient magnitudes and preventing weight norm growth.
6.5 Implementation Dominates Algorithmic Innovation
Engstrom et al. (2020) find that PPO with the clipping mechanism disabled performs indistinguishably from the full algorithm; TRPO (Schulman et al., 2015) with the same code-level additions matches PPO.104104104The non-clipping components that suffice are: observation normalization, reward normalization, value function clipping, global gradient norm clipping, orthogonal weight initialization, and the Adam optimizer. Andrychowicz et al. (2021) identify observation normalization, orthogonal initialization, and learning rate annealing as the three choices accounting for most variance across 250,000 agents and 250 hyperparameter configurations. Huang et al. (2022) catalog 37 implementation details required to reproduce PPO on Atari;105105105These include reward clipping to [1,1][-1,1], frame stacking to 4 frames, a specific episode termination convention, and a numerically stable normalization of the advantage estimate. omitting any produces materially different results.
The most consequential implementation detail is the distinction between termination and truncation (Pardo et al., 2018). In reinforcement learning, an episode can end for two distinct reasons: termination, where the environment reaches a natural absorbing state (the pole falls in CartPole, the robot falls in locomotion tasks), and truncation, where the episode is cut short by an external time limit. At a termination, the value of the successor state is zero: $V(s_{\text{term}})=0$. At a truncation, the episode is merely paused and the successor state has non-zero value: $V(s_{\text{trunc}})\neq 0$. Treating truncated transitions as terminated substitutes zero for a non-zero bootstrap value at every time limit boundary, corrupting every Bellman update in the vicinity of episode boundaries. Pardo et al. (2018) show that this conflation degrades performance by 20–40% on standard MuJoCo benchmarks.106106106The Gymnasium API (Towers et al., 2024) enforces the distinction by returning separate terminated and truncated flags, but most pre-2022 codebases conflate them in the done flag.
6.6 Replay Buffer Pathologies and Reward Scaling
Experience replay (Lin, 1992) decouples the data collection and learning processes, allowing a single transition to be used for multiple gradient updates. The replay ratio (the number of gradient updates per environment step) governs the trade-off between sample efficiency and data staleness. Zhang and Sutton (2017) show that increasing the replay ratio beyond a modest threshold degrades performance.107107107At high replay ratios, the distribution shift between the current policy and the behavior policy that generated the stored data grows large enough to violate the off-policy assumptions of Q-learning. The relationship is non-monotone and environment-dependent, making replay buffer size a sensitive hyperparameter with no universal default.
Schaul et al. (2016) propose prioritized experience replay (PER).108108108PER samples transitions with probability proportional to the magnitude of their TD error, on the argument that high-error transitions are the most informative. Fedus et al. (2020) revisit PER on large-scale Atari experiments and find that uniform sampling from a large enough buffer matches or outperforms PER, while being simpler to implement and tune.
Reward scaling introduces a separate class of failure modes. Standard DQN (Mnih et al., 2015) clips rewards to [1,+1][-1,+1] across all environments to stabilize training. van Hasselt et al. (2016a) observe that reward clipping changes the objective: clipped rewards make all positive events equivalent regardless of magnitude, so the agent learns to maximize the frequency of positive events rather than their cumulative value. This substitution can produce policies that are locally rational under the clipped reward but qualitatively suboptimal under the true reward. van Hasselt et al. (2016a) propose PopArt as a remedy.109109109PopArt normalizes targets to have unit variance while adjusting the output layer so that the policy remains invariant to the normalization. PopArt allows consistent learning across reward scales spanning several orders of magnitude without reward clipping.
Skalse et al. (2022) formalize reward hacking and show that any non-constant proxy reward can in principle be exploited by a sufficiently capable optimizer.110110110Skalse et al. (2022) define a proxy as unhackable if increasing expected proxy return cannot decrease expected true return. Their main result states that for the set of all stochastic policies, two reward functions are unhackable only if one of them is constant. For deterministic policies and finite policy sets, non-trivial unhackable pairs exist, but the conditions are stringent.
6.7 Simulation Study: Bellman Error and Value Error in Offline Policy Evaluation
The MDP uses $s=(k,z)$ with $k$ on a 50-point log-spaced capital grid and $z\in\{0.9,1.1\}$ following a Markov chain with persistence 0.8. Actions are next-period capital choices on the same grid. Reward is $\log(zk^{\alpha}-k^{\prime})$ with $\alpha=0.36$, $\beta=0.96$. Rewards are shifted by $\bar{r}$ (mean reward over feasible pairs) to center $Q^{*}$ near zero, which avoids initialization issues without altering the optimal policy. The offline dataset $\mathcal{D}$ consists of T=2, [^0] T=2{,}000 transitions simulated from the closed-form optimal policy $k^{*}(k,z)=\alpha\beta zk^{\alpha}$. Since the optimal policy concentrates capital near its steady state, $\mathcal{D}$ covers only 11 of the 4,795 feasible $(s,a)$ pairs—0.2% coverage—the distribution mismatch condition in Fujimoto et al. (2022, Corollary 1).
Two algorithms are trained on $\mathcal{D}$.11111111150,000 gradient steps per seed, 3 seeds; two-layer MLP with 64 hidden units, Adam at $5\times 10^{-4}$; target network updated every 500 steps. BRM minimizes $\mathbb{E}_{\mathcal{D}}[(Q(s,a)-(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})))^{2}]$ where both $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ use the current network; gradients flow through both sides simultaneously. The key consequence is that the network can zero the residual at an observed pair $(s,a)$ by co-moving $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ together, rather than moving either toward $Q^{*}$. This is the opposite of supervised learning, where labels are fixed external targets that do not move with the model weights. FQE prevents this by using a frozen target network for $Q(s^{\prime},a^{\prime})$; the network must reduce $Q(s,a)$ toward a target that does not respond to its own gradient steps. Every 500 steps we record the Bellman error on $\mathcal{D}$ (current network on both sides) and the value error $\frac{1}{|\mathcal{F}|}\sum_{(s,a)\in\mathcal{F}}|Q_{\theta}(s,a)-Q^{*}(s,a)|$ over all 4,795 feasible pairs. As a supervised baseline, OLS regression of $\log c$ on $\log k$ and $\log z$ is estimated on expanding windows.112112112Noise $\sigma=0.30$; windows from $n=10$ to 2,0002{,}000; held-out test set of 500 transitions.
The OLS baseline shows tight coupling between training and test loss (Pearson r=1.000r=-1.000; Table 7). The RL results reproduce Fujimoto et al. (2022) in the economic model: BRM achieves Bellman error $816\times$ lower than FQE, yet both methods have nearly identical value error over the full MDP, yielding a VE/BE ratio three orders of magnitude higher for BRM. Both mechanisms from Section 6.1 operate: error cancellation on the 11 observed pairs and unconstrained $Q(s^{\prime},a^{\prime})$ at the 4,784 unobserved pairs.
Method Final BE on $\mathcal{D}$ Final VE (all states) VE/BE ratio
BRM (seed 42) $2.09\times 10^{-4}$ 7.061 33,808
BRM (seed 123) $2.91\times 10^{-4}$ 7.450 25,630
BRM (seed 777) $3.12\times 10^{-4}$ 7.278 23,320
BRM (mean) $2.71\times 10^{-4}$ 7.263 27,586
FQE (seed 42) $0.2037$ 6.927 34
FQE (seed 123) $0.2173$ 7.465 34
FQE (seed 777) $0.2425$ 7.281 30
FQE (mean) $0.2212$ 7.224 33
OLS (n=2,000n=2{,}000) OOS MSE =0.096=0.096 OOS $R^{2}=0.112$ r=1.000r=-1.000
Table 7: Bellman error on dataset $\mathcal{D}$ and value error on the full MDP for BRM and FQE trained on offline Brock–Mirman data. BE is mean squared Bellman error evaluated with the current network on both sides; VE is mean absolute deviation from $Q^{*}$ over all 4,795 feasible state-action pairs.
Figure 12: Left: OLS regression of log consumption on log capital and log productivity, estimated on expanding windows from the Brock–Mirman optimal policy. Out-of-sample MSE (left axis, red) and out-of-sample $R^{2}$ (right axis, blue) track each other with Pearson r=1.000r=-1.000. Right: BRM (orange) and FQE (blue) trained on offline Brock–Mirman data $\mathcal{D}$ (T=2, [^0] T=2{,}000 transitions, 0.2% state-space coverage); note the log scale on the $y$-axis. Solid lines show Bellman error on $\mathcal{D}$ (current network both sides); dashed lines show mean absolute value error against $Q^{*}$ over all 4,795 feasible state-action pairs; shaded bands are $\pm 1$ SE over 3 seeds.
6.8 Discussion and Recommendations
Track episode return and policy entropy alongside training loss; entropy collapse and stagnating return are early warning signs of plasticity loss (Section 6.4). Use PPO or SAC as default baselines before implementing custom algorithms. Report at least 10 seeds per configuration with IQM-based comparisons (Section 6.2).